Abu compares package postpaid mobile plans between Remaja Hebat Plan and Unlimited Plan. Remaja Hebat Plan charges RM0.05 per minutes plus a basic monthly fee for RM 7.50. Unlimited Plan charges RM 0.075 per minute plus a basic monthly fee of RM4.25.
a. Write the model for package postpaid mobile plans for Remaja Hebat Plan and Unlimited Plan.
b. Identify the best mobile plan.
c. Sketch the graph for package postpaid mobile plans between Remaja Hebat Plan and Unlimited Plan

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52778)

.
Abu compares package postpaid mobile plans between Remaja Hebat Plan and Unlimited Plan.
Remaja Hebat Plan charges RM0.05 per minutes plus a basic monthly fee for RM 7.50.
Unlimited Plan charges RM 0.075 per minute plus a basic monthly fee of RM4.25.
a. Write the model for package postpaid mobile plans for Remaja Hebat Plan and Unlimited Plan.
b. Identify the best mobile plan.
c. Sketch the graph for package postpaid mobile plans between Remaja Hebat Plan and Unlimited Plan
~~~~~~~~~~~~~~~~~~~~

(a)  Remaja Hebat Plan is  R(m) = 7.50 + 0.05m    each month, where m represent the called minutes.

     Unlimited    Plan is  U(m) = 4.25 + 0.075m   each month.



(b)  Which plan is the best, it depends on minutes amount, used each month.


     To determine when R-plan is cheaper than U-plan, write and solve inequality

          R(m) < U(M),  or

          7.50 + 0.05m < 4.25 + 0.075m

          7.50 - 4.25 < 0.075m - 0.05m

               3.25   <     0.025m

                   m  > 3.25/0.025 = 130.


     Thus if m > 130 minutes, then R-plan is cheaper than U-plan.

          if m = 130 minutes, both plans are the same.

          if m < 130 minutes, then U-plan is cheaper than R-plan. 



(c)  To make a plot, go to web-site www.desmos.com/calculator/

     Find there free of charge plotting tool for common use.

     Print there expressions for functions and get the plot instantly.

Solved.



Answer by math_tutor2020(3816)


This is a supplement to the previous tutor's answer.
Regarding part (c) and Desmos, you'll need to use x as the input variable instead of m.

Remaja Hebat Plan: y = 0.05x+7.50
Unlimited Plan: y = 0.075x + 4.25

Desmos link
https://www.desmos.com/calculator/xqmtvwaxb2
GeoGebra is another graphing option. If you prefer something like a TI83, then stick to that.